OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 7x+7y=7 x+2y=2

OpenStudy (anonymous):

7x+7y=7 x+2y=2 You want them to be where when you add them, a variable will cancel out so you can solve for the other. multiply top one by 2 14x+14y=14 multiply bottom one by -7 -7x-14y=-14 so you have: 14x+14y=14 -7x-14y=-14 when you add the equations together, the y cancels out. 7x=0 solve for x

OpenStudy (anonymous):

x=0 plug x into one of the original equations to solve for y 0+2y=2 y=1 so your solution would be (0, 1)